### v4

# SCoTER: Structured Chain-of-Thought Transfer for Enhanced Recommendation

Abstract.
Harnessing the reasoning power of Large Language Models (LLMs) for recommender systems is hindered by two fundamental challenges. First, current approaches lack a mechanism for automated, data-driven discovery of effective reasoning patterns, relying instead on brittle manual templates or unstable zero-shot prompting. Second, they employ structure-collapsing integration: direct prompting incurs prohibitive online inference costs, while feature extraction collapses reasoning chains into single vectors, discarding stepwise logic. To address these challenges, we propose SCoTER (Structured Chain-of-Thought Transfer for Enhanced Recommendation), a unified framework that treats pattern discovery and structure-aware transfer as a jointly optimized problem. Specifically, SCoTER operationalizes this through two synergistic components: a GVM pipeline for automated pattern discovery and a structure-preserving integration architecture that transfers stepwise logic to efficient models. Formally, we provide information-theoretic justification proving that structure-preserving transfer achieves tighter performance bounds than structure-agnostic alternatives. Empirically, experiments on four benchmarks demonstrate improvements of 3.75%–11.59% over a strong TIGER backbone. Moreover, in production deployment on the Tencent Advertising Platform, SCoTER achieved a 2.14% lift in Gross Merchandise Value (GMV) while eliminating online LLM inference costs. Overall, SCoTER establishes a principled and production-validated blueprint for transferring structured LLM reasoning to large-scale recommender systems.
Recommender Systems, Large Language Models, Chain-of-Thought, Reasoning Transfer
copyright: acmlicensedjournalyear: 2025isbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systemsccs: Computing methodologies Natural language processingccs: Computing methodologies Knowledge representation and reasoning
1. Introduction
Large Language Models (LLMs) (Zhao et al., 2023; Huang and Chang, 2023) have demonstrated strong reasoning capabilities, especially with Chain-of-Thought (CoT) prompting (Wei et al., 2022). However, transferring this power from objective, logic-driven tasks to the subjective domain of recommender systems introduces two fundamental and interdependent challenges. The first challenge is determining what to transfer. Defining and discovering effective reasoning patterns is inherently difficult without a clear ground truth. User intents are highly diverse, meaning no single pattern can generalize, and sparse data for long-tail items further complicates the validation of any proposed reasoning path. This reveals an inherent conflict between two objectives: preserving the stepwise CoT structure, the source of its reasoning power, and satisfying the stringent low-latency demands of production environments. Addressing these intertwined problems requires a framework that jointly addresses reasoning discovery and structure-preserving transfer.
Figure 1. An overview of the SCoTER framework. It consists of two main components. First, the CoT Pattern Discovery pipeline addresses what to transfer by automatically discovering effective reasoning patterns from data. Second, the Structured CoT Integration architecture addresses how to transfer by integrating these patterns into an efficient model while preserving their step-wise structure.
Despite the clear need for a unified framework, no such solution currently exists. Current approaches address these challenges in isolation, which can be doubly flawed. First, the standalone solutions are inherently inadequate. Pattern discovery for the ’what to transfer’ problem has largely been guided by heuristic, handcrafted schemas conceived from prior assumptions rather than empirical results. This practice decouples pattern design from the downstream recommendation task, resulting in brittle patterns that fail to generalize (Wang et al., 2023a; Zhang et al., ; Wang et al., 2023b). Meanwhile, transfer mechanisms for the ’how to transfer’ problem have typically sacrificed the structural integrity of reasoning, treating it as static features divorced from generative logic (Wang et al., 2024b; Xi et al., 2024; Yang et al., 2025b; Xia et al., 2025). Second, and more critically, this very separation prevents any possibility of joint optimization. It fosters a problematic cycle where patterns are designed without regard for their integration costs, while integration strategies are developed without preserving the patterns’ core logical effectiveness.
To break this problematic cycle, we introduce SCoTER (Structured Chain-of-Thought Transfer for Enhanced Recommendation), a unified framework designed to jointly optimize pattern discovery and structure-preserving integration (Figure 1). To address the ‘what to transfer’ challenge, SCoTER features the Generate-Validate-Mine (GVM) pipeline. This pipeline transforms pattern discovery from a heuristic exercise into a data-driven optimization process: an LLM first generates a diverse set of candidate reasoning paths, which are then validated based on the empirical quality of their recommendations, before a final mining process distills the most effective and generalizable pattern.
For the ‘how to transfer’ challenge, SCoTER employs a lightweight, structure-preserving architecture. This component integrates pre-computed offline reasoning embeddings via an order-aware fusion mechanism, which preserves the sequential structure of CoT while eliminating prohibitive online LLM inference costs. We validate this approach both theoretically and empirically. We provide information-theoretic proof that structure-preserving transfer achieves tighter performance bounds than structure-agnostic alternatives. Furthermore, extensive experiments demonstrate improvements of 3.75-11.59% across four public benchmarks and a 2.14% GMV lift in production deployment.
Our main contributions are:
  • Reasoning Transfer Framework: We establish a systematic framework that unifies pattern discovery and structure preservation as a joint optimization problem, with information-theoretic analysis proving structure-preserving transfer achieves tighter performance bounds.
  • Automated Discovery Pipeline: We introduce the GVM pipeline, replacing manual templates with data-driven selection through latent pattern abstraction.
  • Structure-Preserving Integration: We propose a lightweight architecture using pre-computed stepwise embeddings and order-aware fusion, eliminating online LLM inference while preserving sequential dependencies.
  • Comprehensive Validation: Experiments demonstrate improvements of 3.75-11.59% across four benchmarks and a 2.14% GMV lift in production.
2. Related works
LLM Reasoning for Recommendation. Recent approaches integrating LLM into recommendation systems have explored a variety of complex reasoning structures. CoT-Rec (Liu et al., 2025) employs two-stage prompting for user preference analysis, GOT4Rec (Long et al., 2024) uses Graph-of-Thought frameworks, and ThinkRec (Yu et al., 2025) shifts to System 2 thinking through reasoning data synthesis, while RecGPT (Yi et al., 2025) works to unify multi-step reasoning frameworks. Complementing these efforts, a parallel line of research focuses on refining or distilling reasoning capabilities. This includes the inference-time autoregressive refinement in ReaRec (Tang et al., 2025), the distillation of step-by-step rationales to smaller models by RDRec (Wang et al., 2024a), and the iterative feedback framework used by TrackRec (Xia et al., 2025). However, these methods are limited by relying on heuristic reasoning paths instead of mining user sequences, and their failure to jointly optimize pattern discovery and integration.
Automated Reasoning Discovery: Automated discovery of reasoning patterns has emerged as an alternative to manual template design. Auto-CoT (Zhang et al., ) automatically constructs demonstrations by sampling diverse questions and generating rationales, while Self-prompted CoT (Wang et al., 2023a) enables LLMs to self-induce reasoning steps. Self-Consistency (Wang et al., ) improves reasoning by sampling multiple paths, and broader approaches include APE (Zhou et al., 2022) for automatic prompt engineering, PromptBreeder (Fernando et al., 2024) for evolutionary optimization, and Self-discover (Zhou et al., 2024) for composing atomic reasoning modules. However, these methods are primarily designed for objective tasks with verifiable ground truth. They are less effective in the recommendation domain, which is subjective and has sparse rewards that make it difficult to improve reasoning paths. Our approach, in contrast, addresses this by sampling from broad user behaviors and performing in-depth analysis to use the Recall metric as a dense reward signal.
Figure 2. The proposed Structured Chain-of-Thought Transfer for Enhanced Recommendation (SCoTER) framework. It jointly solves the challenges of what to transfer and how to transfer CoT reasoning. To determine what to transfer, an offline GVM (Generate-Validate-Mine) pipeline automates the discovery of optimal reasoning patterns from data. To address how to transfer these patterns, they are first materialized as step-wise embeddings via Structured Distillation and subsequently fused into a backbone model by a lightweight Online Integration module that preserves the chain’s logical structure.
3. Problem Formulation and Theoretical Foundation
In this section, we establish the theoretical foundation for our framework. We first ground the theory in the context of sequential recommendation by defining core components and our optimization objective. We then provide an information-theoretic justification for our two-pronged approach, namely (i) automated pattern discovery and (ii) structure-preserving integration. We conclude with theoretical guarantees demonstrating that preserving the sequential order of reasoning is provably superior to order-agnostic alternatives.
Formal Definitions We first define the key components of our problem setting.
  • Sequential Recommendation: Given users $\mathcal{U}$ and items $\mathcal{I}$, each user $u\in\mathcal{U}$ has a chronologically ordered interaction history $S=[i_{1},\ldots,i_{T}]\in\mathcal{I}^{T}$. The goal is to learn a model $q_{\theta}$ that approximates the ground-truth next-item distribution $p^{*}(Y|S)$.
  • Reasoning Pattern ($P$): A pattern $P=(p_{1},\ldots,p_{k})\in\mathcal{P}$, with a fixed length $k$, is a high-level reasoning template, e.g., $P=$ (”Analyze history” $\rightarrow$ ”Identify preferences” $\rightarrow$ ”Predict features” $\rightarrow$ ”Recommend items”).
  • Reasoning Chain ($C$): For a sequence $S$ and pattern $P$, a reasoning chain $C=(c_{1},\ldots,c_{k})$ is generated by a pattern-conditioned LLM, denoted $C\sim G_{P}(S)$. Each sentence $c_{j}$ instantiates the template $p_{j}$ with user-specific details. The space of all possible chains is denoted as $\mathcal{C}$.
  • Encoders: We define two types of encoders:
    • An encoder $\psi:\mathcal{C}\to\mathbb{R}^{k\times d}$ is order-sensitive if $\psi(C)\neq\psi(C_{\pi})$ for some permutation $\pi\neq\text{id}$. It represents the chain as a sequence of $k$ step-embeddings (e.g., via Transformers).
    • An encoder $\phi:\mathcal{C}\to\mathbb{R}^{d}$ is order-agnostic if $\phi(C)=\phi(C_{\pi})$ for all permutation $\pi$. It collapses the sequence of step-embeddings into a single $d$-dimensional vector representation (e.g., via mean pooling).
  • $(\rho,\delta)$-Order Sensitivity: A task is $(\rho,\delta)$-order sensitive if with probability at least $\rho$, for a user sequence $S$, a reasoning chain $C$ can be generated whose predictive distribution changes by at least $\delta$ (in TV distance) under step permutation. Formally, $\Pr(S\in\Omega_{\delta})\geq\rho$, where $\Omega_{\delta}=\{S\mid\exists C\sim G_{P}(S),\pi\neq\text{id}\text{ s.t. }\text{TV}(q_{\theta}(\cdot|S,C),q_{\theta}(\cdot|S,C_{\pi}))\geq\delta\}$.
Optimization Objective To jointly identify an optimal pattern $P^{*}$ and train a model $\theta$ that approximates $p^{*}(Y|S)$, our framework maximizes the expected log-likelihood by marginalizing over chains $C\sim G_{P^{*}}(S)$:
(1) $\max_{\theta}\mathbb{E}_{S,Y\sim p^{*}}\left[\log\mathbb{E}_{C\sim G_{P^{*}}(S)}\left[q_{\theta}(Y|S,C)\right]\right]$
This objective effectively decouples pattern discovery (finding $P^{*}$) from model training (optimizing $\theta$).
Information-Theoretic Justification Our framework’s architecture is motivated by decomposing the predictive value of a reasoning chain, $I(C;Y|S)$. Using an operator $f(C)=P$ to extract the pattern from a chain, this value decomposes as:
(2) I(C;Y|S)=I(f(C);Y|S)+I(C;Y|S,f(C))I(C;Y|S)\;=\;I(f(C);Y|S)+I(C;Y|S,f(C))
This decomposition defines the objectives for our two components:
  • Pattern Discovery: The first term, $I(P;Y|S)$, quantifies the pattern’s predictive value. Our GVM pipeline is designed to discover $P^{*}=\arg\max_{P\in\mathcal{P}}I(P;Y|S)$.
  • Structure Preservation: The second term, $I(C;Y|S,P)$, quantifies the value of the chain’s ordered details. Our Structured Integration architecture is designed to preserve this information.
Advantage of Preserving Order We now formalize these advantages through the following results.
Theorem 3.1 (Information-Theoretic Advantage).
Let $\mathcal{H}_{\text{seq}}=\psi(C)$ and $\mathcal{H}_{\text{bag}}=\phi(C)$ be representations from order-sensitive and order-agnostic encoders, respectively. Since $\mathcal{H}_{\text{bag}}$ can be derived from $\mathcal{H}_{\text{seq}}$, the Data Processing Inequality implies:
$I(\mathcal{H}_{\text{seq}};Y|S)\ \geq\ I(\mathcal{H}_{\text{bag}};Y|S)$
Lemma 3.2 (Performance Lower Bound).
For any model $q_{\theta}$, the expected recall is lower-bounded by:
$\mathbb{E}[\operatorname{Recall@K}]\geq\mathbb{E}[m_{K}(S,C)]-\mathbb{E}\left[\mathrm{TV}\left(p^{*}_{S},q_{\theta}(\cdot|S,\mathrm{Encoder}(C))\right)\right]$
where $m_{K}(S,C)$ is the sum of probabilities for the top-$K$ predicted items, and $p^{*}_{S}$ denotes the ground-truth distribution $p^{*}(\cdot|S)$.
Theorem 3.3 (Order-Aware Performance Advantage).
For a $(\rho,\delta)$-order sensitive task, an order-sensitive encoder $\psi$ achieves a performance advantage over an order-agnostic encoder $\phi$:
$\mathbb{E}_{\psi}[\text{Recall@K}]-\mathbb{E}_{\phi}[\text{Recall@K}]$ $\geq\big(\mathbb{E}[m_{K}]_{\psi}-\mathbb{E}[m_{K}]_{\phi}\big)$
$\quad+\frac{\rho\delta}{2}-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\psi})]$
Collectively, these results provide the theoretical guarantee that our structure-preserving approach is provably superior. Detailed proofs are deferred to the Appendix A.
4. Method
Our framework addresses what to transfer by maximizing the mutual-information criterion $I(P;Y|S)$ to obtain $P^{*}$. Specifically, the Generate-Validate-Mine (GVM) pipeline automates this search by generating candidate chains, validating them, and mining the top pattern.
The subsequent challenge is how to transfer this pattern, preserving its step-wise logical structure and retaining the information captured by $I(C;Y|S,P)$. This avoids the structure-collapsing problem exhibited by prior feature extraction methods. To achieve this without the prohibitive cost of online LLM inference, we use a two-stage architecture: the optimal pattern is first distilled into structured representations, which a lightweight fusion module then integrates with the backbone model at serving time.
Figure 2 provides a complete overview of the framework. The following sections are organized around its two core challenges: first, addressing what to transfer using the GVM pipeline, and second, addressing how to transfer it through structured distillation and online fusion.
4.1. What to Transfer: Automated Discovery of Reasoning patterns
Our approach replaces manual template design with a three-phase optimization pipeline: GVM. This process systematically mines the optimal pattern from a diverse set of candidate reasoning chains and extracts it as a symbolic template for the subsequent transfer.
Generate: The Generate phase produces a diverse set of candidate reasoning chains for each user sequence, S. We employ an LLM, such as DeepSeek-R1 (Guo et al., 2025), with a structured prompt (Figure 4) that instructs the model to act as a ”recommendation expert” and defines a specific output format.
The prompt uses a multi-part structure with three distinct outputs: (1) a concise, step-wise reasoning chain in ¡cot_\_path¿ tags that captures the core logic; (2) a detailed elaboration of this logic in a ¡reason¿ block; and (3) a list of 20 ranked recommendations in ¡recommendations¿ tags. This explicit separation is crucial, as it facilitates the subsequent Mine phase by decoupling the abstract reasoning pattern from its detailed explanation.
Two mechanisms are employed to ensure the diversity of the candidate set. First, during generation, we use temperature and top-p nucleus sampling to encourage varied reasoning styles. Second, post-generation, we prune near-duplicate paths using a cosine similarity threshold, $\gamma$. This filtering step preserves semantic diversity and mitigates the over-representation of similar reasoning chains.
Validate: The Validate phase provides a quantitative score for each generated reasoning chain based on its recommendation quality, which serves as the empirical basis for subsequent mining.
This evaluation is formalized using Recall@20, a standard metric for top-K recommendation quality. For each candidate reasoning chain $C$, we compare its list of 20 recommendations, $\widehat{Y}_{20}(C)$, against the ground-truth set of target items, $Y^{*}$. The chain’s performance on a single instance is calculated as:
$\operatorname{Recall@20}(C)\;=\;\frac{|\widehat{Y}_{20}(C)\cap Y^{*}|}{|Y^{*}|}.$
To assess generalized quality, we define Score(C) as the expected Recall@20 across the user distribution. This score measures how consistently a chain produces high-quality recommendations.
$\operatorname{Score}(C)\;=\;\mathbb{E}_{S}\!\big[\operatorname{Recall@20}(C)\big].$
These scores provide an empirical estimate of a chain’s predictive value, allowing the Mine phase to identify patterns that maximize $I(P;Y|S)$ as formalized in Section 3.
Mine: The Mine phase abstracts a single, optimal reasoning pattern from the candidate reasoning chains. Our analysis shows that while the universe of individual chains is intractably large, the underlying pattern space is manageable.
The mining process begins by transforming textual reasoning chains into a dense embedding space using a pre-trained sentence encoder (e.g., Qwen3-8B-Embedding (Zhang et al., 2025b)). Within this space, we perform unsupervised clustering to group semantically similar chains, forming a set of initial candidate patterns.
To select the best pattern, we evaluate these candidates. Our primary criterion is Quality, defined as a pattern’s average effectiveness. For a candidate pattern $P$, we let $\mathcal{C}_{P}$ denote its assigned set of chains. Consistent with the Validate phase, Quality is the mean Recall@20 score across all chains within the cluster:
$\text{Quality}(P)=\frac{1}{|\mathcal{C}_{P}|}\sum\limits_{C\in\mathcal{C}_{P}}\mathrm{Score}(C).$
While average quality is paramount, the final selection also considers Structural Coherence (high intra-pattern semantic similarity) and Performance Stability (low intra-pattern variance in scores). We ultimately choose the pattern that exhibits the best overall balance of these three factors for template extraction.
We conclude the Mine phase by identifying the optimal pattern $P^{*}$ and extracting it as a symbolic, generalizable template. This transformation both provides a human-interpretable artifact for qualitative analysis and serves as a robust instruction set for the subsequent Structured Distillation phase. We achieve this abstraction through a two-stage, LLM-driven synthesis process.
First, we select the top-NN (e.g., $N=10$) chains with the highest cosine similarity to the pattern’s semantic centroid. These exemplars serve as a reliable basis for abstraction. Second, we compile these exemplars into a meta-prompt that directs a powerful LLM to synthesize the shared logical structure. The process culminates in an Optimal CoT Template that captures the core reasoning logic of the discovered pattern.
4.2. How to Transfer: Structure-Preserving Integration
To transfer the discovered pattern $P^{*}$ without structural loss, we employ a two-stage process. First, offline Structured Distillation (Section 4.2.1) materializes the pattern into step-wise embeddings. Second, online Order-Preserving Fusion (Section  4.2.2) integrates these embeddings with the backbone model while preserving sequential dependencies.
4.2.1. Structured Distillation
This stage aims to preserve step-wise structural information and thereby retaining the information captured by $I(C;Y|S,P)$. We achieve this through a structured teacher-student distillation framework.
We leverage the optimal template to guide a powerful teacher LLM (e.g., DeepSeek-R1) in generating structured reasoning chains. For each user sequence $S$ in our training corpus, the teacher model produces template-consistent reasoning $C=(c_{1},c_{2},\ldots,c_{K})$, creating training pairs $\{(S_{i},C_{i})\}_{i=1}^{N}$ where the student learns to generate structured reasoning given user sequences as input.
A smaller, more efficient student model (Qwen3-8B (Yang et al., 2025a)) is fine-tuned on this synthetic dataset, enabling it to generate pattern-consistent reasoning chains that adapt to specific user contexts.
We apply the distilled student model to generate reasoning chains for all data splits. For each sequence $S_{i}$, we feed it through the fine-tuned student model to produce a corresponding reasoning chain $C_{i}=(c_{i,1},c_{i,2},\ldots,c_{i,K})$.
For each generated reasoning step $c_{i,j}$, we extract a dense embedding using a pre-trained sentence encoder (e.g., Qwen3-8B-Embedding (Zhang et al., 2025b)). This process transforms textual reasoning steps into fixed-dimensional embeddings, where $\mathbf{e}_{i,j}\in\mathbb{R}^{D}$ represents the embedding for the $j$-th reasoning step of sequence $S_{i}$:
$\mathbf{e}_{i,j}=\text{SentenceEncoder}(c_{i,j}),\quad j=1,2,\ldots,K.$
The step-wise embeddings for each sequence are then assembled into a structured representation matrix $\mathbf{H}_{i}\in\mathbb{R}^{K\times D}$, which preserves the sequential structure of the reasoning steps:
$\mathbf{H}_{i}=[\mathbf{e}_{i,1};\mathbf{e}_{i,2};\ldots;\mathbf{e}_{i,K}].$
All structured embedding matrices $\{\mathbf{H}_{i}\}_{i=1}^{N}$ are computed and stored offline, enabling the lightweight online fusion phase. This allows for rapid retrieval and integration of pre-computed reasoning representations without incurring generation latency.
4.2.2. Order-Preserving Fusion
This stage integrates pre-computed step-wise embeddings with backbone recommendation models using a lightweight, model-agnostic fusion architecture. This online component prioritizes serving efficiency while preserving the sequential structure critical for reasoning effectiveness.
For each user sequence, we retrieve its corresponding reasoning matrix, $\mathbf{H}_{i}\in\mathbb{R}^{K\times D}$, from the offline repository during inference. An adapter module then projects these reasoning embeddings into the target model’s representation space:
(3) $\mathbf{z}_{i,j}=\text{LayerNorm}(\mathbf{W}_{\text{proj}}\mathbf{e}_{i,j}+\mathbf{b}_{\text{proj}})$
where $\mathbf{e}_{i,j}\in\mathbb{R}^{D}$ is the $j$-th step embedding from $\mathbf{H}_{i}$, $\mathbf{W}_{\text{proj}}\in\mathbb{R}^{d_{\text{item}}\times D}$ projects to the backbone’s item embedding dimension, and $\mathbf{z}_{i,j}\in\mathbb{R}^{d_{\text{item}}}$ is the adapted representation.
To preserve the sequential dependencies critical for structured reasoning, we augment each projected embedding with learnable positional encodings:
(4) $\mathbf{z}_{i,j}^{\text{pos}}=\mathbf{z}_{i,j}+\mathbf{P}_{j}$
where $\mathbf{P}j\in\mathbb{R}^{d{\text{item}}}$ are position embeddings that encode each step’s role within the reasoning sequence.
We employ cross-attention to allow each sequence position to selectively attend to relevant reasoning steps. Let $\mathbf{e}_{\text{seq}}=[\mathbf{e}_{1},\mathbf{e}_{2},\ldots,\mathbf{e}_{T}]$ $\in\mathbb{R}^{T\times d_{\text{item}}}$ denote the backbone model’s embeddings for the user sequence, and $\mathbf{Z}^{\text{pos}}=[\mathbf{z}_{1}^{\text{pos}},\ldots,\mathbf{z}_{K}^{\text{pos}}]\in\mathbb{R}^{K\times d_{\text{item}}}$ represent the projected CoT embeddings with positional encoding. In this cross-attention, the sequence embeddings serve as queries, while the reasoning steps act as both keys and values. The cross-attention mechanism computes attended reasoning representations for each sequence position:
(5) $\mathbf{A}=\text{Attention}(\mathbf{e}_{\text{seq}},\mathbf{Z}^{\text{pos}},\mathbf{Z}^{\text{pos}})$
The attention output is then integrated with the original sequence using adaptive gating:
(6) $\mathbf{g}$ $=\sigma(\mathbf{W}_{g}[\mathbf{e}_{\text{seq}};\mathbf{A}]+\mathbf{b}_{g})$
(7) $\mathbf{E}_{\text{fused}}$ $=\text{LayerNorm}(\mathbf{g}\odot\mathbf{e}_{\text{seq}}+(1-\mathbf{g})\odot\mathbf{A})$
Here, $\mathbf{e}_{\text{seq}};\mathbf{A}]$ represents concatenation along the feature dimension, and the final layer normalization is applied to the gated output.
To align the reasoning space with the recommendation objective, we employ a contrastive learning component using the InfoNCE loss (Oord et al., 2018). The loss is computed between the final reasoning step embedding, $\mathbf{z}_{K}$, and the target item embedding, $\mathbf{v}_{\text{target}}$:
(8) $\mathcal{L}_{\text{InfoNCE}}=-\log\frac{\exp(\text{sim}(\mathbf{z}_{K},\mathbf{v}_{\text{target}})/\tau)}{\sum_{j=1}^{B}\exp(\text{sim}(\mathbf{z}_{K},\mathbf{v}_{j})/\tau)}$
The term $\text{sim}(\cdot,\cdot)$ represents cosine similarity, $\tau$ is the temperature parameter, and $B$ is the batch size, with $\{\mathbf{v}_{j}\}_{j=1}^{B}$ including the target item and negative samples from other batch items. The full training objective combines the recommendation loss with the contrastive alignment loss:
(9) $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{rec}}+\lambda\mathcal{L}_{\text{InfoNCE}}$
where $\lambda$ is a hyperparameter that controls the contribution of the contrastive term.
This structured integration architecture preserves the step-wise nature of CoT reasoning, allowing downstream models to leverage both the progressive reasoning flow and the final recommendation-oriented representations to improve prediction accuracy.
Table 1. Performance comparison across traditional and generative recommendation methods. Our SCoTER consistently outperforms existing baselines. Best results in bold, second-best underlined.
Baseline Methods Our Approach
Dataset Metric MF LightGCN Caser HGN Bert4Rec SASRec TIGER SCoTER Improve vs TIGER
Beauty Recall@5 0.0202 0.0228 0.0279 0.0344 0.0203 0.0387 0.0392 0.0434 10.71%
Recall@10 0.0379 0.0421 0.0456 0.0564 0.0347 0.0605 0.0594 0.0656 10.44%
NDCG@5 0.0122 0.0136 0.0172 0.0214 0.0124 0.0249 0.0257 0.0276 7.39%
NDCG@10 0.0178 0.0198 0.0229 0.0284 0.0137 0.0318 0.0321 0.0347 8.10%
Instruments Recall@5 0.0738 0.0757 0.0770 0.0813 0.0671 0.0857 0.0865 0.0908 4.97%
Recall@10 0.0967 0.1010 0.0995 0.1048 0.0822 0.1083 0.1062 0.1110 4.52%
NDCG@5 0.0473 0.0472 0.0639 0.0668 0.0560 0.0715 0.0736 0.0765 3.94%
NDCG@10 0.0547 0.0554 0.0711 0.0774 0.0608 0.0788 0.0799 0.0829 3.75%
Sports Recall@5 0.0087 0.0098 0.0116 0.0189 0.0115 0.0233 0.0233 0.0260 11.59%
Recall@10 0.0165 0.0184 0.0194 0.0313 0.0191 0.0350 0.0379 0.0406 7.12%
NDCG@5 0.0053 0.0061 0.0072 0.0120 0.0075 0.0154 0.0150 0.0161 7.33%
NDCG@10 0.0079 0.0087 0.0097 0.0159 0.0099 0.0192 0.0197 0.0209 6.09%
Yelp Recall@5 0.0220 0.0248 0.0150 0.0186 0.0186 0.0183 0.0241 0.0258 7.05%
Recall@10 0.0381 0.0403 0.0263 0.0326 0.0291 0.0296 0.0385 0.0406 5.45%
NDCG@5 0.0138 0.0156 0.0099 0.0115 0.0115 0.0116 0.0158 0.0174 10.13%
NDCG@10 0.0190 0.0207 0.0134 0.0159 0.0159 0.0152 0.0204 0.0222 8.82%
5. Experiments
In this section, we conduct a series of experiments to answer the following research questions:
RQ1: How does our proposed SCoTER framework perform against the sequential and generative recommendation models?
RQ2 (What to transfer): How effective is our automated reasoning pattern discovery compared to manual, heuristic-based CoT templates?
RQ3 (How to transfer): How do structure-preserving components contribute to reasoning transfer effectiveness?
RQ4: Does integration with a backbone model synergize collaborative and reasoning signals more effectively than standalone LLM generation?
Table 2. Dataset statistics of the evaluation benchmarks. “AvgLen” represents the average length of item sequences.
Dataset #Users #Items #Interactions AvgLen
Beauty 22,363 12,101 198,502 8.88
Instruments 24,772 9,922 206,153 8.32
Sports 35,598 18,357 296,337 8.32
Yelp 30,431 20,033 316,354 10.40
5.1. Experimental Setup
Datasets. We conduct experiments on four widely used datasets: three subsets of the Amazon Product Reviews dataset (He and McAuley, 2016; Ni et al., 2019) (Beauty, Instruments, and Sports) and the Yelp dataset. Table 2 summarizes the statistics of these datasets. Following previous work (Zhou et al., 2020), we process the data to enforce a 5-core density, removing all users and items with fewer than five interactions. All user sequences are then normalized to a uniform length of 20 through padding or truncation, preserving their most recent interactions. For evaluation, we use the leave-one-out protocol: each user’s final interaction is designated for testing, the penultimate one for validation, and the remaining interactions are used for training
Baselines. We compare our proposed method, Structured Chain-of-Thought Recommendation (SCoTER), against a comprehensive suite of representative baselines that span different paradigms:
  • MF (Koren et al., 2009): A classic model that uses matrix factorization to learn latent embeddings for users and items.
  • LightGCN (He et al., 2020): A graph convolutional network that captures collaborative signals via neighborhood aggregation.
  • Caser (Tang and Wang, 2018): A sequential model that employs convolutional neural networks to capture local sequential patterns.
  • HGN (Ma et al., 2019): A sequential model that utilizes a hierarchical gating network to adaptively integrate a user’s long- and short-term preferences.
  • SASRec (Kang and McAuley, 2018): A sequential model that uses a self-attention mechanism to capture long-range dependencies and dynamic user preferences.
  • Bert4Rec (Sun et al., 2019): A sequential model that uses a deep bidirectional self-attention mechanism to model user sequences.
  • TIGER (Rajput et al., 2023): A generative model that represents items as discrete token sequences, enabling recommendation through autoregressive decoding. We select TIGER as the backbone for our method due to its strong generative performance and architectural compatibility with reasoning integration.
  • SCoTER: Structured-CoT enhances the TIGER backbone by integrating structured Chain-of-Thought reasoning, as detailed in Section 4.
Evaluation Protocol. Performance is evaluated using two standard top-K ranking metrics: Recall@K and NDCG@K. Following common practice, we report the main results for $K\in\{5,10\}$. To ensure a fair evaluation and avoid sampling bias, we perform a full ranking over the entire item catalog for each user.
Implementation Details. For traditional methods, we follow standard implementations with hyperparameters tuned on validation sets. For generative methods, we adopted a unified configuration based on the T5 architecture. The backbone is a 4-layer Transformer, configured with a model dimension of 128, six attention heads (dimension 64), a 1024-unit hidden MLP, ReLU activation, and a 0.1 dropout rate. The tokenizer employs RQ-VAE for discrete semantic encoding with 4 codebooks, each containing 256 embeddings of dimension 32. Semantic inputs to RQ-VAE are derived from the embeddings of item titles and descriptions processed by Qwen3-8B-Embedding (Zhang et al., 2025b). During inference, we use a beam size of 20 to balance recommendation quality and efficiency.
For SCoTER, which enhances the TIGER backbone, we apply multi-head cross-attention (6 heads) between the sequence embeddings $\mathbf{e}_{\text{seq}}$ and pre-computed offline reasoning embeddings $\mathbf{Z}^{\text{pos}}$. We use learnable positional embeddings to preserve sequential dependencies and adaptive gating with sigmoid activation to control the fusion of sequence and reasoning representations. Training uses Adam optimizer with learning rate $2\times 10^{-4}$, weight decay $5\times 10^{-5}$, and 200 epochs with early stopping. The contrastive learning weight $\lambda$ is set to 0.1.
Table 3. Ablation results on Beauty dataset.
Variant Recall@5 Recall@10 NDCG@5 NDCG@10
Full model 0.0434 ()(-) 0.0656 ()(-) 0.0276 ()(-) 0.0347 ()(-)
w/o Position 0.0424 ($\downarrow 2.30\%$) 0.0647 ($\downarrow 1.37\%$) 0.0270 ($\downarrow 2.17\%$) 0.0341 ($\downarrow 1.73\%$)
w/o Contrastive 0.0413 ($\downarrow 4.84\%$) 0.0639 ($\downarrow 2.59\%$) 0.0267 ($\downarrow 3.26\%$) 0.0337 ($\downarrow 2.88\%$)
w/o Step-wise CoT embedding 0.0407 ($\downarrow 6.22\%$) 0.0624 ($\downarrow 4.88\%$) 0.0265 ($\downarrow 3.99\%$) 0.0335 ($\downarrow 3.46\%$)
Tiger 0.0392 ($\downarrow 9.68\%$) 0.0594 ($\downarrow 9.45\%$) 0.0257 ($\downarrow 6.88\%$) 0.0321 ($\downarrow 7.49\%$)
5.2. Performance Comparison (RQ1)
To assess the overall effectiveness of our reasoning transfer framework, we compare SCoTER against seven strong baseline models. The comprehensive results are presented in Table 1.
Our method consistently outperforms all baseline models across every dataset and metric. It achieves significant performance gains over TIGER, ranging from 3.75% to 11.59%. The most substantial improvements are observed on the Beauty and Sports datasets, highlighting the effectiveness of our method in diverse domains. Notably, the uplift is often more pronounced in top-5 metrics (Recall@5, NDCG@5) compared to top-10 metrics, suggesting that structured reasoning particularly benefits precision-critical scenarios where top recommendations must be accurate.
Among the baselines, SASRec stands out as the top-performing traditional method, while TIGER demonstrates strong generative capability. Despite its strengths, the model lacks systematic mechanisms for reasoning pattern optimization and fails to preserve order-aware reasoning representations. The ability of our method to elevate TIGER’s performance demonstrates that incorporating explicit reasoning pattern discovery and structure-aware integration addresses these limitations and provides substantial value.
These empirical results validate the effectiveness of our overall framework design. The consistent gains across diverse datasets and metrics demonstrate that SCoTER successfully transfers reasoning capabilities to enhance recommendation performance.
5.3. Automated Pattern Discovery (RQ2)
To rigorously evaluate our automated discovery, we compared the GVM-discovered pattern against several manual templates. These manual templates, detailed in the Appendix B, represent general-purpose reasoning structures derived from domain knowledge and expert intuition. As shown in Figure 3, when integrated with the TIGER backbone, the GVM-discovered pattern demonstrates a substantial advantage over the manual alternatives across all metrics on the Beauty dataset. Notably, its 10.71% improvement in Recall@5 nearly doubles the gain of the best-performing manual template, establishing a robust performance gap with gains ranging from 7.39% (NDCG@5) to 10.71%.
This superiority extends beyond integrated settings to standalone LLM generation (Table 4). On both a fine-tuned Qwen3-8B and the larger DeepSeek-R1, our pattern consistently outperforms the manual alternative. This consistent outperformance across diverse settings and models highlights a fundamental architectural advantage, which stems from the GVM pipeline’s systematic, data-driven approach.
This advantage can be understood by deconstructing the GVM process. Manual templates usually rely on generalized human experience. While providing a reasonable starting point, this generality prevents them from capturing the fine-grained, dynamic signals specific to current user interactions, thus limiting their practical effectiveness. Conversely, our GVM pipeline systematically uncovers superior patterns. The Generate phase explores a vast landscape of potential reasoning patterns directly from the data, moving beyond pre-defined assumptions. Crucially, the Validate phase acts as an empirical filter, scoring each candidate based on its actual recommendation performance, thereby creating a feedback loop that ensures only data-supported reasoning paths survive. Finally, the Mine phase distills the most effective and generalizable logic from this validated set. This systematic discovery process allows us to identify latent, data-specific reasoning structures that are not just theoretically sound, but empirically proven to be more beneficial.
5.4. Structure-Preserving Integration (RQ3)
To validate our structure-preserving architecture, we conducted a systematic ablation study (Table 3). The results demonstrate that each component is essential, with their removal causing measurable performance degradation ranging from 2.30% to 6.22% in Recall@5.
Step-wise CoT embedding emerges as the most critical component, yielding the largest performance degradation, with the Recall@5 score dropping by 6.22%. It preserves the progressive refinement inherent in reasoning chains. Each step builds upon previous insights to iteratively narrow the recommendation space. Collapsing this multi-step structure into a single vector discards these intermediate logical dependencies, forcing the model to recommend without the benefit of stepwise deliberation.
Table 4. LLM-as-recommender performance on Beauty dataset. Models generate recommendations directly from reasoning chains without backbone integration.
DeepSeek-R1 Qwen3-8B (Fine-tuned)
Recall@20 NDCG@20 Recall@20 NDCG@20
Two-step 0.0078 0.0041 0.0340 0.0138
Three-step 0.0089 0.0047 0.0344 0.0142
Five-step 0.0098 0.0052 0.0352 0.0145
SCoTER 0.0105 0.0056 0.0363 0.0152
Beyond this structural foundation, positional encoding and contrastive learning provide complementary enhancements. First, positional encoding preserves sequential order. Without explicit positional signals, the model struggles to differentiate between an early hypothesis exploration and a final refinement. This ambiguity hinders the application of appropriate attention weights across different reasoning stages, thereby degrading the model’s ability to leverage the sequential structure. Consequently, removing this component leads to a significant 2.30% drop in Recall@5. Second, contrastive learning aligns this reasoning with recommendation objectives. It provides a crucial supervisory signal that steers the logic beyond mere internal coherence to match user preferences. Its removal, therefore, causes an even larger degradation, with the Recall@5 score dropping by 4.84% to 0.0413.
Finally, the study reveals a synergistic effect that amplifies their individual contributions. Removing both positional encoding and contrastive learning simultaneously results in a performance drop greater than the sum of their individual impacts. This indicates a cooperative relationship: positional encoding preserves the sequential logic, while contrastive learning aligns this logic with recommendation objectives.
Figure 3. Performance Improvement Relative to Tiger on Beauty with backbone integration. Manual CoT Templates (Two-step, Three-step, Five-step) are compared with the automatically Discovered CoT.
$5.$ Integration Synergy (RQ4)
A pivotal insight is revealed when comparing the outcomes of standalone LLM-based recommendations against our fully integrated model. The best direct-generation configuration—a fine-tuned Qwen3-8B using our CoT pattern—achieves a Recall@20 of 0.0363. In contrast, the integrated approach reaches a substantially higher Recall@10 of 0.0656. This gap highlights the fundamental value of fusing complementary information sources.
Our architecture’s advantage stems from its ability to synergize two distinct modalities. LLM generation relies on explicit semantic logic but lacks the implicit collaborative signals that are the foundation of modern recommenders, such as latent patterns of item co-occurrence or user taste clusters. To bridge this gap, the recommender backbone provides strong collaborative priors, while the CoT module injects an interpretable reasoning layer. This fusion creates recommendations that are both empirically grounded and logically justified—a capability neither component possesses alone.
Beyond synergy, the results reveal another key insight: task-specific adaptation is more critical than raw model scale. This is demonstrated by the smaller, fine-tuned Qwen3-8B consistently outperforming the much larger DeepSeek-R1. This outcome validates our structured distillation, demonstrating its ability to transfer sophisticated reasoning into an efficient model. Ultimately, this confirms a viable path for integrating LLM reasoning into large-scale, production-ready systems.
Table 5. Relative improvement of our online A/B testing on the Tencent Advertising Platform.
Online Metric Relative Lift
GMV (Overall) +2.14%
GMV (Sparse Users) +4.10%
GMV (Dense Users) +1.49%
Negative Feedback Rate -0.24%
”Not Interested” Rate -0.25%
5.6. Online A/B Test
We validated the real-world effectiveness of SCOTER through deployment on the Tencent Advertising Platform. Grounded in promising offline results (a +6.1% relative lift in HitR@100 metrics), we initiated an online A/B test. Using a 5% traffic experimental group, we compared SCOTER against our online model for one week, with Gross Merchandise Value (GMV) as the primary metric.
As reported in Table 5, SCOTER delivered a significant +2.14% lift in overall GMV. Furthermore, a stratified analysis revealed that the performance gains were most pronounced for users with sparse interaction histories, achieving a +4.1% GMV lift. This contrasts with the +1.49% lift for users with dense histories, highlighting its significant potential to mitigate the data sparsity problem.
SCOTER also demonstrated positive trends in user experience. We observed a 0.24% decrease in the average negative feedback rate and a 0.25% decrease in the ”not interested” rate. These results indicate that the recommendations generated by SCOTER are not only more profitable but also better aligned with user preferences.
We refer interested readers to (Zhang et al., 2025a) for a more comprehensive discussion of the online experimental setup, additional metrics, and extended analysis.
6. Conclusion
In this paper, we identify and address two challenges in applying CoT reasoning to recommendation: discovering effective reasoning patterns beyond brittle and hand-crafted heuristics, and transferring them to efficient models without collapsing their essential stepwise logic under low-latency demands. To tackle these challenges in a unified manner, we propose SCoTER, a novel framework featuring an automated GVM pipeline for pattern discovery and a structure-preserving architecture. The efficacy of this framework is validated on both theoretical and empirical grounds. In principle, our analysis establishes the advantage of preserving reasoning structure. In practice, comprehensive experiments demonstrate that SCoTER not only consistently outperforms state-of-the-art baselines but also achieves a 2.14% lift in production GMV. Together, these results establish SCoTER as a systematic and empirically-grounded methodology for integrating structured LLM reasoning into recommender systems.
Appendix A Complete Theoretical Analysis
A.1. Formal Definitions
Definition A.1.
  • Permutation Operation: Let $\pi\in\mathfrak{S}_{k}$ be a non-identity permutation. For a reasoning chain $C=(c_{1},\ldots,c_{k})$, the permuted chain is $C_{\pi}=(c_{\pi(1)},\ldots,c_{\pi(k)})$.
  • Total Variation (TV) Distance: For distributions $p$ and $q$, the distance is $\text{TV}(p,q)=\frac{1}{2}\sum_{y}|p(y)-q(y)|$.
  • Order-Sensitive Set: $\Omega_{\delta}=\{S:\exists\pi\neq\text{id},\text{TV}(p^{*}(\cdot|S,C),$ $p^{*}(\cdot|S,C_{\pi}))\geq\delta\}$, with probability mass $\rho=\mathbb{P}(S\in\Omega_{\delta})$.
  • Encoder Classification: An encoder $\phi$ is order-agnostic if $\phi(C)=\phi(C_{\pi})$ for all $\pi$. An encoder $\psi$ that does not satisfy this is order-sensitive.
A.2. Information-Theoretic Advantage
Theorem A.2 (Information-Theoretic Advantage).
Let $\mathcal{H}_{seq}=\psi(C)$ and $\mathcal{H}_{bag}=\phi(C)$ be representations from order-sensitive and order-agnostic encoders, respectively. Then:
$I(\mathcal{H}_{seq};Y|S)\geq I(\mathcal{H}_{bag};Y|S)$
Proof.
Since the order-agnostic representation $\mathcal{H}_{bag}$ is a deterministic function of the order-sensitive representation $\mathcal{H}_{seq}$ (e.g., via mean pooling), the result follows directly from the Data Processing Inequality, which states that no processing of a variable can increase its mutual information with another variable. ∎
A.3. Performance Lower Bound
Theorem A.3 (Basic Performance Lower Bound).
For any model $q_{\theta}$, the expected recall satisfies:
$\mathbb{E}[\text{Recall@K}]\geq\mathbb{E}[m_{K}]-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\theta})]$
where $m_{K}$ is the model’s probability mass for its top-K predictions.
Proof.
Step 1: For a given sequence $S$, the expected recall is:
$\mathbb{E}[\text{Recall@K}|S]=\mathbb{E}[\mathbf{1}_{Y\in A_{q}}|S]=p^{*}_{S}(A_{q})$
where $A_{q}$ is the set of top-K predicted items.
Step 2: By the definition of TV distance, for any event $A_{q}$:
$|p^{*}_{S}(A_{q})-q_{\theta}(A_{q}|S)|\leq\text{TV}(p^{*}_{S},q_{\theta})$
Step 3: From Step 2, we have:
$p^{*}_{S}(A_{q})\geq q_{\theta}(A_{q}|S)-\text{TV}(p^{*}_{S},q_{\theta})$
Step 4: By definition:
$q_{\theta}(A_{q}|S)=\sum_{y\in A_{q}}q_{\theta}(y|S)=m_{K}$
Step 5: Combining Steps 1, 3, and 4:
$\mathbb{E}[\text{Recall@K}|S]\geq m_{K}-\text{TV}(p^{*}_{S},q_{\theta})$
Step 6: Taking the expectation over all samples $S$:
$\mathbb{E}[\text{Recall@K}]\geq\mathbb{E}[m_{K}]-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\theta})]$
A.4. Order-Agnostic Encoder’s Fitting Loss
Lemma A.4 (Collision Penalty).
For any $S\in\Omega_{\delta}$, there exists a permutation $\pi\neq\text{id}$ such that for an order-agnostic encoder $\phi$:
$\max\{\text{TV}(p^{*}(\cdot|S,C),q_{\phi}(\cdot|\phi(C))),\text{TV}(p^{*}(\cdot|S,C_{\pi}),q_{\phi}(\cdot|\phi(C)))\}\geq\frac{\delta}{2}$
Proof.
Since $S\in\Omega_{\delta}$, we have $\text{TV}(p^{*}(\cdot|S,C),p^{*}(\cdot|S,C_{\pi}))\geq\delta$. An order-agnostic encoder $\phi$ produces the same prediction $q_{\phi}$ for both $C$ and $C_{\pi}$. By the triangle inequality:
$\delta$ $\leq\text{TV}\!\big(p^{*}(\cdot|S,C),\,p^{*}(\cdot|S,C_{\pi})\big)$
$\leq\text{TV}\!\big(p^{*}(\cdot|S,C),\,q_{\phi}\big)+\text{TV}\!\big(q_{\phi},\,p^{*}(\cdot|S,C_{\pi})\big)$
If both terms on the right were less than $\delta/2$, their sum would be less than $\delta$, which is a contradiction. Thus, at least one of the terms must be greater than or equal to $\delta/2$. ∎
Lemma A.5 (Expected Fitting Error Lower Bound).
For an order-agnostic encoder $\phi$:
$\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]\geq\rho\cdot\frac{\delta}{2}$
Proof.
We decompose the expectation over sensitive and non-sensitive samples:
$\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]$ $=\rho\cdot\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})\,|\,S\in\Omega_{\delta}]$
$\quad+(1-\rho)\cdot\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})\,|\,S\notin\Omega_{\delta}]$
From the previous lemma, the conditional expectation for sensitive samples is at least $\delta/2$. Since the second term is non-negative, the result follows. ∎
A.5. Main Theorem: Order-Aware Performance Advantage
Theorem A.6 (Order-Aware Performance Advantage).
The performance advantage of an order-sensitive encoder $\psi$ over an order-agnostic encoder $\phi$ is:
$\mathbb{E}_{\psi}[\text{Recall@K}]-\mathbb{E}_{\phi}[\text{Recall@K}]$ $\geq\big(\mathbb{E}[m_{K}]_{\psi}-\mathbb{E}[m_{K}]_{\phi}\big)$
$\quad+\frac{\rho\delta}{2}-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\psi})]$
Proof.
Step 1: We apply the Performance Lower Bound from Theorem A.2 to each encoder:
(10) $\mathbb{E}_{\psi}[\text{Recall@K}]$ $\geq\mathbb{E}[m_{K}]_{\psi}-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\psi})]$
(11) $\mathbb{E}_{\phi}[\text{Recall@K}]$ $\geq\mathbb{E}[m_{K}]_{\phi}-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]$
Step 2 : Subtracting the second inequality from the first gives:
Advantage $\geq(\mathbb{E}[m_{K}]_{\psi}-\mathbb{E}[m_{K}]_{\phi})$
$\quad+\big(\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\psi})]\big)$
Step 3: Using the inherent error bound for the order-agnostic encoder from Lemma A.4, $\mathbb{E}[\text{TV}(p^{*}_{S},q_{\phi})]\geq\rho\cdot\frac{\delta}{2}$, we arrive at the final result:
$\text{Advantage}\geq(\mathbb{E}[m_{K}]_{\psi}-\mathbb{E}[m_{K}]_{\phi})+\frac{\rho\delta}{2}-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\psi})]$
Appendix B Manual CoT Templates
This section details the manual Chain-of-Thought (CoT) templates used for comparison against our GVM-discovered pattern. These templates represent heuristic-based reasoning structures of increasing complexity. For the experiments, the following strings were used to guide the LLM’s reasoning process inside the ‘¡cot_\_path¿‘ tag.
Two-step Template
A direct, two-stage reasoning path focusing on mining interests and then recommending.
<cot_path>User Interest Mining -¿ Item Tag Prediction & Recommendation </cot_path>
Prompt: Inferring Recommendation Paths from User Behavior You are an expert in recommendation algorithms. Based on the user’s historical behavior data, please infer a reasonable analysis and recommendation path, and predict 20 different product features that the user may be interested in the future. Requirements: (1) First output <cot_path>, with a reasoning chain of at least two steps, each step $\leq$8 words, connected with ”-¿”. (2) <reason> must be placed after <cot_path>, and the content must strictly follow each step in <cot_path> to expand the reasoning one by one. (3) Finally output 20 recommended product feature descriptions. Output format requirements: <cot_path> [Summarize the core reasoning path here: Step 1 -¿ Step 2-¿ $\cdots$ -¿ Item recommendation] </cot_path> <reason> [Explain the reasoning logic step by step according to the steps in cot_path] </reason> <recommendations> <item>Detailed feature description of product 1, including category, brand, function, characteristics, etc.</item> <item>Detailed feature description of product 2, including category, brand, function, characteristics, etc.</item> $\cdots$ <item>Detailed feature description of product 20, including category, brand, function, characteristics, etc.</item> </recommendations>
Figure 4. The structure of the prompt used to guide the LLMs for generating candidate CoT paths and recommended item features. The prompt requires a structured output including a concise reasoning path (<cot_path>), detailed reasoning (<reason>), and 20 specific product feature recommendations (<recommendations>).
Three-step Template
This template adds an explicit intermediate step for summarizing a user profile before interest extraction.
<cot_path >User Profile Summary -¿ User Interest Extraction -¿ Item Recommendation </cot_path>
Five-step Template
A more granular template that breaks down the analysis into multiple distinct phases, from data analysis to feature prediction before the final recommendation.
<cot_path >Behavioral Data Analysis -¿ Interest Pattern Recognition -¿ Preference Trend Analysis -¿ Predictive Feature Generation -¿ Targeted Item Recommendation </cot_path>